# Exploring the interplay between mass-energy equivalence, interactions and entanglement in an optical lattice clock

### v1

@@ Abstract
Furthermore, we find that the simultaneous presence of single-atom GRS and photon-mediated interactions can lead to quantum entanglement as shown in Fig. 3(d,e). In fact, in the regime $\omega_{\rm split}\ll\Delta E$, the projection of the wave function into the fully symmetric manifold imposed by the many-body gap transforms the single-particle term into an effective one-axis twisting (OAT) [51, 52] interaction term $\chi\hat{S}^{z}\hat{S}^{z}$, with $\chi\sim\omega_{\rm split}^{2}/[N(\Delta E)]$ (see Ref. [53] where the splitting is generated by a different mechanism). In this case, entanglement builds up for $t>t_{\rm syn}$, as witnessed by a squeezing parameter [52], $\xi^{2}\equiv\min_{\varphi}N(\Delta S_{\varphi}^{\perp})^{2}/|\langle\hat{\mathbf{S}}\rangle|^{2}<1$ (see Fig. 3(d)). Here, $(\Delta S_{\varphi}^{\perp})^{2}$ is the variance of spin noise along an axis perpendicular to the collective spin $\langle\hat{\mathbf{S}}\rangle$. A faster growth of entanglement can be seen in the regime $\omega_{\rm split}\sim\Delta E$ (see Fig. 3(e)). Since the entanglement in this case is not captured by spin squeezing, instead we characterize the entanglement by the normalized Rényi entropy $\tilde{S}_{N/2}=-2\log_{2}(\mathrm{tr}(\rho_{N/2}^{2}))/N$, where $\rho_{N/2}$ is the reduced density matrix by taking partial trace over half of the system. The entanglement builds up for $\omega_{\rm split}t/2\pi>1$ in this case, which might be due to population transfer to highly entangled states in manifolds of lower total spin [42]. For the implementation of entanglement generation, $\omega_{\rm split}/NJ_{\perp}\sim 0.1-1$ is achievable for $10$ cm -- $1$ m separation and $NJ_{\perp}/2\pi\sim 0.1$ Hz.
@@
To study the effects of the GRS on quantum entanglement, we consider the scenario with entangled initial states, such as the ones generated using cavity induced OAT interactions [54, 55], $\hat{U}_{\mathrm{OAT}}(Q)=\exp(-iQ\hat{S}^{z}\hat{S}^{z}/N)$, where $Q$ is the shearing strength (Fig. 4(a)). The squeezing direction corresponds to the direction with minimum value of $(\Delta S_{\varphi}^{\perp})^{2}$, which can be controlled by performing a rotation $\hat{R}_{x}^{\theta}=\exp(-i\theta\hat{S}^{x})$ as shown in Fig. 4. We assume the initial state preparation is fast enough such that the GRS does not play a role. We use $|\psi_{0}\rangle$ to denote the state after OAT interactions, and $|\psi(\theta)\rangle$ for the state after the $\hat{R}_{x}^{\theta}$ rotation.
@@
In Fig. 4(b), we show that it is possible to control $t_{\mathrm{syn}}$ below or above the value of a product initial state $t_{\mathrm{syn},0}$ (obtained in Fig. 3) depending on the rotation $\hat{R}_{x}^{\theta}$. The ratio $t_{\mathrm{syn}}/t_{\mathrm{syn},0}$ under $\hat{H}_{\rm cav}$ ($J_{\perp}=J_{z}$) can be understood using the following analytic result [42],
@@
where ${\rm Cov}(\alpha,\beta)\equiv\langle\psi(\theta)|(\hat{S}^{\alpha}\hat{S}^{\beta}+\hat{S}^{\beta}\hat{S}^{\alpha})|\psi(\theta)\rangle-2\langle\psi(\theta)|\hat{S}^{\alpha}|\psi(\theta)\rangle\langle\psi(\theta)|\hat{S}^{\beta}|\psi(\theta)\rangle$, with $\alpha,\beta=x,y,z$. The tunability of $t_{\mathrm{syn}}$ is due to the $\theta$-dependence of ${\rm Cov}(y,z)$. The tunable range $\Delta t_{\mathrm{syn}}\equiv\max_{\theta}t_{\mathrm{syn}}-\min_{\theta}t_{\mathrm{syn}}$ can be used as a measure of entanglement (see Fig. 4(c)), since $4{\rm Cov}(y,z)$ is approaching the quantum Fisher information $F_{Q}$ [52], which corresponds to the maximal eigenvalue of the matrix $F_{Q,\alpha\beta}=2{\rm Cov}(\alpha,\beta)$.
Conclusion and outlook.—We discussed protocols accessible in OLCs that explore how the single-atom GR effects modify many-body dynamics generated by photon-mediated interactions. A similar interplay should be observable with atomic superexchange interactions [42]. While so far we have mostly focused on highly localized atomic arrays, generalizations to the case of itinerant particles where motion and other GR corrections also become relevant, will open unique opportunities for testing the basic tenets of GR when extended into the complex quantum domain.
@@ Acknowledgements.
We thank James Thompson, Raphael Kaubruegger, Dylan Young, Alexander Aeppli, Michael Werner and Torsten Zache for useful discussions. This work is supported by the Sloan Foundation, the Simons Foundations and the Heising-Simons Foundation, NSF JILA-PFC PHY-2317149 and NSF OMA-2016244 (QLCI) grants, and by NIST.

### v2

@@ Abstract
Furthermore, we find that the simultaneous presence of single-atom GRS and photon-mediated interactions can lead to quantum entanglement as shown in Fig. 3(d,e). In fact, in the regime $\omega_{\rm split}\ll\Delta E$, the projection of the wave function into the fully symmetric manifold imposed by the many-body gap transforms the single-particle term into an effective one-axis twisting (OAT) [51, 52] interaction term $\chi\hat{S}^{z}\hat{S}^{z}$, with $\chi\sim\omega_{\rm split}^{2}/[N(\Delta E)]$ (see Ref. [53] where the splitting is generated by a different mechanism). In this case, entanglement builds up for $t>t_{\rm syn}$, as witnessed by a squeezing parameter [52], $\xi^{2}\equiv\min_{\varphi}N(\Delta S_{\varphi}^{\perp})^{2}/|\langle\hat{\mathbf{S}}\rangle|^{2}<1$ (see Fig. 3(d)). Here, $(\Delta S_{\varphi}^{\perp})^{2}$ is the variance of spin noise along an axis perpendicular to the collective spin $\langle\hat{\mathbf{S}}\rangle$. A faster growth of entanglement can be seen in the regime $\omega_{\rm split}\sim\Delta E$ (see Fig. 3(e)). Since the entanglement in this case is not captured by spin squeezing, instead we characterize the entanglement by the normalized Rényi entropy $\tilde{S}_{N/2}=-2\log_{2}(\mathrm{tr}(\hat{\rho}_{N/2}^{2}))/N$, where $\hat{\rho}_{N/2}$ is the reduced density matrix by taking partial trace over half of the system. The entanglement builds up at a time scale $t_{\rm split}\sim\pi/\omega_{\rm split}$ in this case, which might be due to population transfer to highly entangled states in manifolds of lower total spin. For the implementation of entanglement generation, $\omega_{\rm split}/NJ_{\perp}\sim 0.1-1$ is achievable for $10$ cm -- $1$ m separation and $NJ_{\perp}/2\pi\sim 0.1$ Hz. Therefore, the time scale for entanglement growth is at the order of $10$s and thus requires careful suppression of systematics.
@@
To study the effects of the GRS on quantum entanglement, we consider the scenario with entangled initial states, such as the ones generated using cavity induced OAT interactions [54, 55], $\hat{U}_{\mathrm{OAT}}(Q)=\exp(-iQ\hat{S}^{z}\hat{S}^{z}/N)$, where $Q$ is the shearing strength (Fig. 4(a)). The squeezing direction corresponds to the direction with minimum value of $(\Delta S_{\varphi}^{\perp})^{2}$, which can be controlled by performing a rotation $\hat{R}_{x}^{\theta}=\exp(-i\theta\hat{S}^{x})$ as shown in Fig. 4. As demonstrated in Ref. [55], entangled state preparation time scales in an OLC are at the order of $10$ ms, so it is reasonable to ignore GRS in initial state preparation. We use $|\psi_{0}\rangle$ to denote the state after OAT interactions, and $|\psi(\theta)\rangle$ for the state after the $\hat{R}_{x}^{\theta}$ rotation.
@@
In Fig. 4(b), we show that it is possible to control $t_{\mathrm{syn}}$ below or above the value of a product initial state $t_{\mathrm{syn},0}$ ($Q=0$, as obtained in Fig. 3) depending on the rotation $\hat{R}_{x}^{\theta}$. The ratio $t_{\mathrm{syn}}/t_{\mathrm{syn},0}$ under $\hat{H}_{\rm cGR}$ ($J_{\perp}=J_{z}$) can be understood using the following analytic result [41],
@@
where ${\rm Cov}(\alpha,\beta)\equiv\langle\psi(\theta)|(\hat{S}^{\alpha}\hat{S}^{\beta}+\hat{S}^{\beta}\hat{S}^{\alpha})|\psi(\theta)\rangle-2\langle\psi(\theta)|\hat{S}^{\alpha}|\psi(\theta)\rangle\langle\psi(\theta)|\hat{S}^{\beta}|\psi(\theta)\rangle$, with $\alpha,\beta=x,y,z$. The tunability of $t_{\mathrm{syn}}$ is due to the $\theta$-dependence of ${\rm Cov}(y,z)$. The tunable range $\Delta t_{\mathrm{syn}}\equiv\max_{\theta}t_{\mathrm{syn}}-\min_{\theta}t_{\mathrm{syn}}$ can be used as a measure of entanglement (see Fig. 4(c)), since $C_{yz}=4\max_{\theta}{\rm Cov}(y,z)$ is approaching the quantum Fisher information $F_{Q}$ [52], which corresponds to the maximal eigenvalue of the matrix $F_{Q,\alpha\beta}=2{\rm Cov}(\alpha,\beta)$.
Conclusion and outlook.—We discussed protocols accessible in OLCs that explore how the single-atom GR effects modify many-body dynamics generated by photon-mediated interactions. A similar interplay should be observable with atomic superexchange interactions. While so far we have mostly focused on highly localized atomic arrays, generalizations to the case of itinerant particles where motion and other GR corrections also become relevant, will open unique opportunities for testing the basic tenets of GR when extended into the complex quantum domain.
@@ Acknowledgements.
We thank James Thompson, Raphael Kaubruegger, Dylan Young, Alexander Aeppli, Michael Werner and Torsten Zache for useful discussions. This work is supported by the Sloan Foundation, the Simons Foundation and the Heising-Simons Foundation, NSF JILA-PFC PHY-2317149 and NSF OMA-2016244 (QLCI) grants, and by NIST. VML and KH acknowledge funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through Project-ID 274200144 – SFB 1227 (project A09) and Project-ID 390837967 - EXC 2123.
Appendix A Appendix: Derivations of the dressing scheme
Construction of dressed states.—Here we explain the dressing scheme in Fig. 2 in more detail. Due to magnetic Zeeman shifts, the clock transitions between different nuclear spins are frequency resolved, thus we can restrict the dynamics within three internal levels $|e,m_{F}\rangle$, $|g,m_{F}\rangle$ and $|g,m_{F}-1\rangle$. A dressing laser (with Rabi frequency $\Omega$, laser frequency $\omega_{d}$) is used to couple $|e,m_{F}\rangle$ with $|g,m_{F}-1\rangle$ states (see Fig. 2(a)), leading to the following Hamiltonian,
$\hat{H}_{\rm dress}/\hbar$ $=\omega_{0}\hat{\mathcal{P}}_{e,m_{F}}-\omega_{Z}\hat{\mathcal{P}}_{g,m_{F}-1}$ (7)
$+\frac{\Omega}{2}(|e,m_{F}\rangle\langle g,m_{F}-1|e^{-i\omega_{d}t}+h.c.),$
where $\omega_{Z}$ is the Zeeman shift between $|g,m_{F}\rangle$ and $|g,m_{F}-1\rangle$ states, and we set the energy of $|g,m_{F}\rangle$ state to 00. The projection operators are defined by $\hat{\mathcal{P}}_{\psi}=|\psi\rangle\langle\psi|$. We then perform a unitary transformation $\hat{U}=\exp(-i\omega_{d}t\,\hat{\mathcal{P}}_{g,m_{F}-1})$ to rewrite the Hamiltonian in the rotating frame of the dressing laser,
$\hat{H}_{\rm dress}^{\prime}/\hbar$ $=\omega_{0}\hat{\mathcal{P}}_{e,m_{F}}+(\omega_{0}+\delta)\hat{\mathcal{P}}_{g,m_{F}-1}$ (8)
$+\frac{\Omega}{2}(|e,m_{F}\rangle\langle g,m_{F}-1|+h.c.),$
where $\delta=\omega_{d}-\omega_{0}-\omega_{Z}$ is the detuning of the dressing laser. The eigenstates of this Hamiltonian are given by $|+\rangle=C_{1}|e,m_{F}\rangle+C_{2}|g,m_{F}-1\rangle$, $|-\rangle=-C_{2}|e,m_{F}\rangle+C_{1}|g,m_{F}-1\rangle$ with energy $E_{\pm}/\hbar=\omega_{0}+\delta/2\pm\sqrt{\Omega^{2}+\delta^{2}}/2$. These are the dressed states defined in the main text with coefficients $C_{1}=(1-\delta/\sqrt{\Omega^{2}+\delta^{2}})^{1/2}/\sqrt{2}$ and $C_{2}=(1+\delta/\sqrt{\Omega^{2}+\delta^{2}})^{1/2}/\sqrt{2}$.
Tunability of GR effects.—Firstly we analyze the effects of the dressing scheme on the GR effects. We focus on the GRS described by the following Hamiltonian,
$\hat{H}_{\rm GRS}=(\Delta M_{0})\phi\hat{\mathcal{P}}_{e,m_{F}}=$ $\,(\Delta M_{0})\phi\bigg{(}|C_{2}|^{2}\hat{\mathcal{P}}_{-}+|C_{1}|^{2}\hat{\mathcal{P}}_{+}$ (9)
$+\Big{(}C_{1}C_{2}^{*}|+\rangle\langle-|\,+h.c.\Big{)}\bigg{)},$
where $\Delta M_{0}=\hbar\omega_{0}/c^{2}$ is the mass defect without the dressing protocol. As shown in Fig. 2(a), we probe the transition between $|g,m_{F}\rangle$ and |ket|$\rangle$| - ⟩. Consider the case that the energy gap $\sqrt{\Omega^{2}+\delta^{2}}$ between |+ket|$\rangle$| + ⟩ and |ket|$\rangle$| - ⟩ states due to $\hat{H}_{\rm dress}$ is much larger than $\omega_{0}\phi/c^{2}$, the coupling between |+ket|$\rangle$| + ⟩ and |ket|$\rangle$| - ⟩ states in $\hat{H}_{\rm GRS}$ are suppressed significantly. Therefore, to the leading order, when projected into the relevant $|g,m_{F}\rangle$ and |ket|$\rangle$| - ⟩ states we can write $\hat{H}_{\mathrm{GRS}}$ as:
$\hat{H}_{\mathrm{GRS}}\approx(\Delta M_{0})\phi|C_{2}|^{2}\hat{\mathcal{P}}_{-}=(\Delta M)\phi\hat{\mathcal{P}}_{-},$ (10)
which is equivalent to a modification of the mass defect (see Fig. 2(b)),
$\Delta M=|C_{2}|^{2}\Delta M_{0}.$ (11)
The same analysis applies to the SDS.
The tunability of gravitational redshift due to the dressing scheme allows us to distinguish GRS from a magnetic field gradient. In the dressing scheme, $\hat{H}_{\rm GRS}$ leads to a position-dependent correction of the dressed state energy $E_{-}$,
$E_{-}(Z)=E_{-}+\frac{\hbar\omega_{0}g_{\rm LO}}{c^{2}}|C_{2}|^{2}Z,$ (12)
which can be resolved via clock spectroscopy in the effective two level system formed by $|g,m_{F}\rangle$ and |ket|$\rangle$| - ⟩ states. While for a small magnetic field gradient term adding on top of a constant magnetic field, we have $\omega_{0}(Z)=\omega_{0}+(\eta_{e}-\eta_{g})m_{F}Z$, $\omega_{Z}(Z)=\omega_{Z}+\eta_{g}Z$, where $\eta_{e}=-\mathcal{G}_{{}^{3}P_{0}}\mu_{B}\partial_{Z}B$, $\eta_{g}=-\mathcal{G}_{{}^{1}S_{0}}\mu_{B}\partial_{Z}B$, with $\mathcal{G}_{{}^{3}P_{0}}$ and $\mathcal{G}_{{}^{1}S_{0}}$ representing the Landé g-factors, and $\mu_{B}$ is the Bohr magneton. So the position-dependent correction due to magnetic field gradients is given by
$E_{-}(Z)=E_{-}+\bigg{[}|C_{2}|^{2}(\eta_{e}-\eta_{g})m_{F}-|C_{1}|^{2}\eta_{g}\bigg{]}Z.$ (13)
Since $\eta_{g}\neq 0$, we find different dependence by varying $\delta$ compared to the GRS (see Eq. (12)). The reason is that different ground-state nuclear spins have the same mass but different Zeeman shifts.
Figure 5: Schematic of the cavity-mediated interactions in the dressed basis. The cavity axis is along $Z$ and the quantization axis is along magnetic field $B$ labeled in the plot. The additional nuclear spin state in the dressing protocol allows for coupling to two polarization modes in the cavity and enhancing the tunability of the cavity-mediated interactions.
Tunability of cavity-mediated interactions.—Secondly we analyze the effects of the dressing protocol on the cavity-mediated interactions. We focus on multilevel alkaline earth atoms with the quantization axis for nuclear spins perpendicular to the cavity axis (see Fig. 5). In this case, the two polarization modes supported by the cavity can drive the $\pi$ transition and the linear combination of $\sigma^{+}$ and $\sigma^{-}$ transitions, so we can define the multilevel raising operators for these two polarization modes, $\hat{\Pi}^{+}=\sum_{jm}C^{0}_{m}|e,m\rangle_{j}\langle g,m|$, $\hat{\Sigma}^{+}=\sum_{jm}\frac{i}{\sqrt{2}}(C^{-1}_{m}|e,m-1\rangle_{j}\langle g,m|+C^{+1}_{m}|e,m+1\rangle_{j}\langle g,m|)$, where $j$ is the label of atoms, $m$ is the label of nuclear spins, and $C^{\sigma}_{m}\equiv\langle F,m;1,\sigma|F,m+\sigma\rangle$ are the Clebsch-Gordan coefficients. The photon-mediated exchange interactions for multilevel alkaline earth atoms take the following form [61, 62],
$\hat{H}_{c}/\hbar=\chi(\hat{\Pi}^{+}\hat{\Pi}^{-}+\hat{\Sigma}^{+}\hat{\Sigma}^{-}),$ (14)
[…]
$\hat{H}_{c}/\hbar\approx J_{\perp}\hat{S}^{+}\hat{S}^{-}+J_{z}\bigg{(}\frac{N}{2}+\hat{S}^{z}\bigg{)}^{2},$ (15)
where
$J_{\perp}=\chi(C^{0}_{m_{F}})^{2}|C_{2}|^{2},\quad J_{z}=\chi\frac{(C^{+1}_{m_{F}-1})^{2}}{2}|C_{1}|^{2}|C_{2}|^{2}.$ (16)
The tunability of $J_{\perp}$ and $J_{z}$ is shown in Fig. 2(c). Eq. (5) describing the interplay between gravitational redshift and cavity-mediated interactions is a combination of Eq. (10) and Eq. (15) in the Wannier-Stark basis.
The physical meaning of the $J_{\perp}$ term is the spin exchange interactions between $|g,m_{F}\rangle$ and |ket|$\rangle$| - ⟩ state via the $\hat{\Pi}^{+}\hat{\Pi}^{-}$ process, i.e. an atom in |ket|$\rangle$| - ⟩ state emits a virtual photon into the cavity and flips to $|g,m_{F}\rangle$ state ($\hat{\Pi}^{-}$), and another atom in $|g,m_{F}\rangle$ state absorbs the same photon and flips to |ket|$\rangle$| - ⟩ state ($\hat{\Pi}^{+}$). The physical meaning of $J_{z}$ term is the collective frequency shift of |ket|$\rangle$| - ⟩ state via the $\hat{\Sigma}^{+}\hat{\Sigma}^{-}$ process, i.e. an atom in |ket|$\rangle$| - ⟩ state absorbs a photon from the dressing laser beam and emits it into the cavity while staying in the same state ($\hat{\Sigma}^{-}$), then another atom in |ket|$\rangle$| - ⟩ state absorbs the same cavity photon and emits it back to the dressing laser beam ($\hat{\Sigma}^{+}$).
Experimental considerations.—Finally we focus on the experimental requirements of the dressing scheme. For 87Sr atoms, we consider the Zeeman shifts between nuclear spin states to be at the order of $10^{2}$ Hz. In order to frequency resolve a single transition between nuclear spin states, we have $\Omega/2\pi,\delta/2\pi\sim 10$ Hz. Considering $\chi N/2\pi$ at the order of Hz for clock transition as shown in Ref. [44], and GRS at the order of mHz for $1$ cm spatial separation, the validity of the dressing scheme is ensured.
As for experimental implementation, spatial inhomogeneities exist for the atomic frequency and the parameters $\delta$ and $\Omega$ in $E_{-}$. The effects of GRS will be washed out if the inhomogeneities are much larger than the value of GRS. For $1$ cm spatial separation, the mHz scale of GRS requires to control the inhomogeneities below $10^{-4}$ Hz for direct observation.
For the atomic frequencies and detuning $\delta$, the leading order contributions are from spatial inhomogeneties in the magnetic field. One can suppress first order Zeeman shifts by probing opposite nuclear spin states and calculating the averaged frequency, which is attainable based on the parameters in Ref. [7]. This approach allows for cancellation of first order Zeeman shifts up to shot-to-shot fluctuations, and residue effects of the magnetic field can be distinguish from GRS based on the dressing scheme.
For the dressing laser Rabi frequency $\Omega$, the leading order contributions are from the spatial profile of the laser beam. If we denote the modification of $\Omega$ as $\Delta\Omega$, the change of $E_{\pm}$ is given by
$\Delta E_{\pm}/\hbar=\pm\frac{1}{2}\sqrt{\Omega^{2}+\delta^{2}}\bigg{(}1+\frac{(\Delta\Omega)\Omega}{\Omega^{2}+\delta^{2}}\bigg{)}.$ (17)
The suppression of inhomogeneties in $E_{-}$ requires $\Delta\Omega/2\pi<10^{-4}~{}$z, which is equivalent to $\Delta\Omega/\Omega<10^{-5}$. In principle, this requirement is achievable using an ultrastable cavity, which allows for precise control of the spatial mode of the dressing laser. Notice that $\Delta E_{+}+\Delta E_{-}=0$, an alternative approach to reduce this requirement is to average the transition frequency of $|g,m_{F}\rangle\leftrightarrow|-\rangle$ and $|g,m_{F}\rangle\leftrightarrow|+\rangle$,
$\frac{E_{-}(Z)+E_{+}(Z)}{2}=\hbar\omega_{0}+\frac{\hbar\delta}{2}+\frac{1}{2}\frac{\hbar\omega_{0}}{c^{2}}\phi(Z).$ (18)
In this way, the averaged transition frequency becomes independent of $\Omega$. Even though one sacrifices the full tunability of GRS, it is still changed to half of its value without dressing.